Carly owns stock in a company that consistently paid a growing dividend over the last five years. The first year Carly owned the stock, she received $1.71 per share and in the fifth year, she received $2.89 per share. What is the growth rate of dividends during this time?

7%

14%

12%

5%
Answer #1
FV= PV*(1+r)^n
Where,
FV= Future Value
PV = Present Value
r = Interest rate
n= periods in number
2.89= $1.71*( 1+r)^4
2.89/1.71 = (1+r)^4
1.69006 =(1+r)^4
r= 0.14018
Growth rate = 14%